How do you solve the following equation? ln X + ln(x+3)= 1
Thanks for using your time to help me out.

Answer by stanbon(75887)

solve the following equation? ln X + ln(x+3)= 1
--------------------------------------------------
Comment: These ln terms are exponents.
When you are adding them you are multiplying.
----------------------------------------------
ln(x(x+3)) = 1
x(x+3) = e^1
x^2 + 3x - e = 0
--------------
Note: e = 2.71828....
----------------------
Use the quadratic formula to solve for "x":
x = [-3 +- sqrt(9 - 4*1*-e)]/2
x = [-3 +- sqrt(9+4e)]/2
x = [-3 +- 4.4579]/2
---
Positive solution:
x = 0.7290..
Cheers,
Stan H.